# SeNMo: A Self-Normalizing Deep Learning Model for Enhanced Multi-Omics Data Analysis in Oncology

### v1

@@ 2.3 SeNMo Deep Learning Model
In learning scenarios with hundreds to thousands of features and relatively few training samples, Feedforward networks are susceptible to overfitting [102]. Unlike Convolutional Neural Networks (CNNs), the weights in Feedforward networks are shared, making them more prone to training instabilities caused by perturbations and regularization techniques like stochastic gradient descent and Dropout. CNNs do not perform as well on high-dimensional, low-sample data for several reasons including the spatial invariance assumption, fixed input size, and parameter efficiency viz-a-viz multiomics data sparsity. Transformers-based models are also not inherently optimized for high-dimensional, low-sample data, such as in genomics or other multiomics datasets, due to the fact that these models use attention mechanism to predict the next token in case of language tasks and meaningful patterns in non-language tasks, which fails in case of highly sparse molecular data. To address overfitting in high-dimensional, low-sample-size multiomics data and employ more robust regularization techniques during training, we draw inspiration from Self-Normalizing Networks (SNN) introduced by Klambauer et al. [105]. SNNs have been extensively used for their ability to effectively handle high-dimensional data with limited samples, making them particularly relevant in the field of multiomics data analysis. Our learning framework is based on the stacked layers of SNNs, as described below.
@@
As shown in Figure 4, SeNMo comprises of stacked blocks of SNN layers, where each block is composed of a linear unit, SELU activation, followed by Alpha dropout. Combined, these blocks enable high-level abstract representations by keeping neuron activations converged towards zero mean and unit variance [105]. Linear unit is essentially equivalent to what is commonly referred to as a “fully connected (FC) layer" or a “multilayer perceptron (MLP) layer" in traditional neural network architectures. SELU activations are an alternative to the traditional rectified linear unit (ReLU) activations commonly used in neural networks. Klambauer et al. demonstrated through the Banach fixed-point theorem that activations with close proximity to zero mean and unit variance, propagating through numerous network layers, will ultimately converge to zero mean and unit variance [105]. SELU activations have the unique property of self-normalization, meaning that the activations tend to converge to a mean of zero and a standard deviation of one, regardless of the input distribution. The mathematical equation for the SELU activation function is:
$selu(x)=\lambda\begin{cases}x&\text{if }x>0\\ \alpha(e^{x}-1)&\text{if }x\leq 0\end{cases}$ (5)
where:
  • $\lambda$ is a scaling factor (typically set to $1.05071$).
@@
  • $\alpha$ is the negative scale factor (typically set to $1.6733$).
@@
Dropout is a regularization technique that randomly sets a fraction of input units to zero during training to prevent overfitting. Alpha dropout is a modified version of traditional dropout regularization, specifically designed to preserve the self-normalizing property of SELU activations. Alpha Dropout applies a dropout mask to the activations during training, scaled by a factor that ensures the mean and variance of the activations remain unchanged. This scaling factor is computed based on the dropout rate and the SELU parameters ($\lambda$ and $\alpha$). Mathematically,
$\text{AlphaDropout}(x)=\frac{x-\mu(x)}{std(x)}\times mask+\mu(x)$ (6)
where:
  • $x$ is the input activation.
  • $\mu(x)$, $std(x)$ are mean and standard deviation of the input activation, respectively.
@@
  • $mask$ is a binary mask generated with the specified dropout rate.
@@
Together, SELU activations and Alpha Dropout enable SeNMo blocks to maintain a stable mean and variance of activations throughout the network layers, facilitating more stable training and better generalization performance. Similarly, SELU activation and alpha dropout help mitigate training instabilities caused by vanishing or exploding gradients in the Feedforward networks. Our network architecture consists of seven fully-connected hidden layers followed by Exponential Linear Unit (ELU) activation and alpha dropout, where the number of neurons in each block is shown in the inset of Figure 4. The final fully-connected layer is utilized to learn a latent representation of the sample, which we call the patient embedding $x\in\mathbbm{R}^{48}$.
Figure 4: Architecture of the SeNMo encoder network. There are seven hidden layers each comprising of a linear unit, SELU activation, and Alpha dropout. The trained model has 83.33 million parameters. The number of neurons in each hidden layer, input layer and output layer are also depicted in the figure. The same model is used for regression and classification tasks.
@@ 2.4.1 Evaluation
We evaluate SeNMo’s performance with the quantitative and statistical metrics common for survival outcome prediction and grade classification. For survival analysis, we evaluate the model using the Concordance Index (C-Index). For the primary cancer type classification, we generate the classification report comprising of average accuracy, average precision, recall, F1-score, confusion matrix, and scatter plot. For testing the statistical significance of the predictions, we use the Log Rank Test. Below we explain the evaluation metrics and statistical test in detail.
@@
    Loss Function: The loss being used for backpropagation in the model is a combination of three components: cox loss, cross-entropy loss, and regularization loss. This combined loss function aims to simultaneously optimize the model’s ability to predict survival outcomes (Cox loss), encourage model simplicity or sparsity (regularization loss), and model the likelihood of cancer types (cross-entropy loss). The overall loss is a weighted sum of these three components, where each component is multiplied by a corresponding regularization hyperparameter ($\lambda_{c}$, $\lambda_{ce}$, $\lambda_{r}$). This weighted sum allows for balancing the influence of each loss component on the optimization process. Mathematically, the overall loss can be expressed as:
@@
    • Cox proportional hazards loss ($L_{cox}$): The Cox loss is a measure of dissimilarity between the predicted hazard scores and the true event times in survival analysis. It is calculated using the Cox proportional hazards model and penalizes deviations between predicted and observed survival outcomes of all individuals who are at risk at time $t_{i}$, weighted by the censoring indicator [106]. The function takes a vector of survival times for each individual in the batch, censoring status for each individual (1 if the event occurred, 0 if censored), and the predicted log hazard ratio for each individual from the neural network, and returns the Cox loss for the batch, which is used to train the neural network via backpropagation. This backpropagation encourages the model to assign higher hazards to high-risk individuals and lower predicted hazards to censored individuals or who experience the event later. Mathematically, the Cox loss is expressed as:
@@
      where:
      • $N$ is the batch size (number of samples).
      • $\theta_{i}$ is the predicted hazard for sample $i$.
      • $R_{ij}$ is the indicator function that equals 1 if the survival time of sample $j$ is greater than or equal to the survival time of sample $i$, and 0 otherwise.
      • $censor_{i}$ is the censoring indicator for sample $i$, which equals 1 if the event is observed for sample $i$ and 0 otherwise.
    • Cross-entropy loss ($L_{ce}$): The cross-entropy loss is a common loss function used for multi-class classification problems, particularly when each sample belongs to one of $C$ classes. When combined with a LogSoftmax layer, the function measures how well a model’s predicted log probabilities match the true distribution across various classes. For a multi-class classification problem having $C$ classes, the model’s outputs (raw class scores or logits) are transformed into log probabilities using a LogSoftmax layer. The cross-entropy loss compares these log probabilities to the true distribution, which is usually represented in a one-hot encoded format. The loss is calculated by negating the log probability of the true class across all samples in a batch and then averaging these values. For the given output of LogSoftmax, $\log(p_{n,c})$ for each class $c$ in each sample $n$, the cross-entropy loss for a multi-class problem can be defined as:
@@
      where:
      • $N$ is the total number of samples.
      • $C$ are the total classes.
      • $y_{n,c}$ is the target label for sample $n$ and class $c$, typically 1 for the true class and 0 otherwise.
@@
    Concordance Index (C-Index): The C-Index is a frequently used evaluation metric in survival analysis to assess the predictive accuracy of a model for the time-to-event outcomes [10]. It measures the degree to which the model’s predictions correlate with the actual survival times observed in the data. It quantifies the model’s ability to correctly rank pairs of subjects based on their predicted survival times. The C-Index evaluates the probability that, in a randomly selected pair of individuals, the one who experienced the event (like death or failure) first also had a higher risk score predicted by the model. Risk score is the output of the survival model and represents the expected order of the events; the higher the score, the higher the risk of experiencing the event sooner [10]. We used the $concordance\_index$ Lifelines function to calculate the C-Index. This function takes the predicted hazard scores for each individual, the true event indicators (e.g., 1 if an event occurred, 0 if censored) for each individual, and the survival times (time to event or censoring) for each individual. The C-Index function computes the fraction of all pairs of subjects whose predicted event times are correctly ordered among all pairs where one subject experienced an event and the other did not. C-Index ranges between 0 and 1 where 0.5 is the expected result from random predictions, 1.0 is a perfect concordance and, 0.0 is perfect anti-concordance [10]. Mathematically,
@@
    where:
    • Concordant pairs are pairs of individuals where the predicted survival times are correctly ordered relative to the observed survival times.
    • Tied pairs are number of pairs where the predictions are equal or survival times are the same.
    • Total number of evaluable pairs are the total pairs considered, excluding pairs with censoring issues or other exclusions.
    • $\hat{S}_{i}$ and $\hat{S}_{j}$ represent the predicted risks or survival probabilities for individuals $i$ and $j$, respectively.
    • $T_{i}<T_{j}$ implies that individual $i$ experienced the event before individual $j$.
    • $\delta_{i}=1$ indicates that the event for individual $i$ was observed (not censored).
@@
    Cox log-rank function: The Cox Log-rank function calculates the p-value using the log-rank test based on predicted hazard scores, censor values, and the true overall survival times. The log-rank test is a statistical method to compare the survival distributions of two groups, where the null hypothesis is that there is no difference between the groups. It is commonly used in survival analysis to compare the observed number of events in each group to the number of events expected under the null hypothesis. For the hazard ratio $h_{i}(t)$ of group $i$ at time $t$ is given by,
    $\begin{split}&H_{0}:h_{1}(t)=h_{2}(t)\\ &H_{A}:h_{1}(t)=ch_{2}(t),\;\;c\neq 1\end{split}$ (11)
@@
    where:
    • $O_{i}$ is the observed number of events at time point $i$ in the sample.
    • $E_{i}$ is the expected number of events at time point $i$ under the null hypothesis.
    • $N$ is the total number of observed time points.
2.4.2 Hyperparameters Search
Hyperparameters are non-learnable parameters of a deep learning model and are crucial as they govern the learning process and model architecture. Hyperparameter tuning involves selecting the optimal combination of parameters that results in the best model performance. Common hyperparameters include learning rate and policy, batch size, number of epochs, weight decay, dropout type and probability, and architecture specifics such as the number of hidden layers and neurons in each layer. Methods for hyperparameter search range from grid search, where all possible combinations of parameters are evaluated; to random search, which randomly samples parameter combinations within predefined bounds. More sophisticated techniques like Bayesian optimization or using automated machine learning (AutoML) tools can dynamically adjust parameters based on previous results to find the best solutions more efficiently. We employed weights and biases [108] utility to carry out random and Bayesian methods of hyperparameters search. The list of hyperparameters we searched for training and fine-tuning is given in Table 3. For model training, we conducted around 400 simulations to find the current hyperparameters. For fine-tuning the SeNMo model on CPTAC-LUSC and Moffitt’s Lung SCC data, we ran around 1000 and 450 simulations, respectively. The plots for these simulations are given in the appendix figures 9, 10, and 11.
Hyperparams Training (range) Finetuning (range)
Learning Rate [1e-6, 1e-1] [1e-8, 1e-3]
Weight Decay [1e-6, 1e-1] [1e-4, 1e-2]
Dropout [0.1, 0.65] [0.05, 0.45]
Batch Size [64, 128, 256, 512] [8, 16, 32, 48]
Epochs [50, 100] [8, 10, 15, 20, 30]
Hidden Layers [1, 2, 3, 4, 5, 6, 7, 8, 9] -
Hidden Neurons [2048, 1024, 512, 256, 128, 48, 32] -
Optimizer [adam, sgd, rmsprop, adamw] [adam, adamw]
Learning Rate Policy [linear, exp, step, plateau, cosine] [linear, exp, plateau]
Frozen Layers - [7, 6, 5, 4, 3, 2]
Additional Layers - [1, 2, 3]
Table 3: Hyperparameters for training and finetuning.
2.4.3 Frameworks, Compute resources, and wall-clock times
We trained SeNMo model using the Moffitt Cancer Center’s HPC machine with 1 Tesla V100 32GB GPU running Ubuntu 22.04.4 and CUDA 12.2. The entire code was developed in Python and PyTorch frameworks. The software frameworks and corresponding packages used in our codebase are given in Table 4. Training time for our current 83.33 Million parameter SeNMo encoder is approximately 11 hours. We conducted the hyperparameters search of the pan-cancer model for approximately 20 days using multiple GPUs in parallel. Finetuning the trained model on a given data having around 150 patients approximately takes 15 minutes. It took us one week to find the fine-tuning hyperparameters for the CPTAC-LSCC data.
@@ 2.4.3 Frameworks, Compute resources, and wall-clock times
Figure 5: Pan-cancer C-Index results for OS prediction. We trained and evaluated the SeNMo model on different combinations of data modalities. Training and validation was carried out on the $80\%$ of the total data, whereas inference was done on the $20\%$ held-out test set. As the number of modalities increased in the pan-cancer data, the model’s performance improved as depicted by the upward trend of C-Index. All the results shown here are statistically significant, i.e., $p<0.05$.

### v2

@@ 2.3 SeNMo Deep Learning Model
In scenarios involving hundreds or thousands of features with relatively few training samples, feedforward networks often face the risk of overfitting [101]. Unlike CNNs, weights in feedforward networks are shared, making them vulnerable to training instabilities caused by perturbations and regularization techniques such as stochastic gradient descent and dropout. CNNs, on the other hand, struggle to handle high-dimensional, low-sample data due to the spatial invariance assumption, fixed input size, and inefficiencies in managing multi-omics data sparsity. Transformer-based models are also suboptimal for high-dimensional, low-sample data, as they rely heavily on attention mechanisms tailored for predicting sequential patterns, which fails when dealing with highly sparse molecular data.
To address the challenges of overfitting and instability in high-dimensional, low-sample-size multi-omics data, we drew inspiration from self-normalizing networks introduced by Klambauer et al. [106]. Self-normalizing neural networks are particularly suited for high-dimensional datasets with limited samples, a characteristic that makes them highly relevant for multi-omics analysis. The SeNMo architecture is based on stacked layers of self-normalizing neural networks, as detailed below.
@@
As illustrated in Figure 4, SeNMo comprises stacked blocks of self-normalizing neural network layers, where each block includes a linear unit, a Scaled Exponential Linear Unit (SELU) activation, and Alpha-Dropout. These components enable high-level abstract representations while keeping neuron activations close to zero mean and unit variance [106]. The linear unit is equivalent to a "fully connected" or MLP layer commonly used in traditional neural network architectures. Klambauer et al. demonstrated through the Banach fixed-point theorem that activations with close proximity to zero mean and unit variance, propagating through numerous network layers, will ultimately converge to zero mean and unit variance [106]. SELU activations, an alternative to traditional rectified linear unit activations, offer a self-normalizing effect, ensuring activations converge to zero mean and unit variance regardless of the input distribution. The SELU activation function is expressed mathematically as:
$\text{SELU}(x)=\lambda\begin{cases}x&\text{if }x>0\\ \alpha(e^{x}-1)&\text{if }x\leq 0\end{cases}$ (5)
@@
where, $\lambda$ is a scaling factor (typically set to $1.05071$) and $\alpha$ is the negative scale factor (typically set to $1.6733$).
@@
Dropout, a regularization method that randomly sets a fraction of input units to zero during training, prevents overfitting. Alpha-Dropout, a modified version of traditional dropout, is designed to maintain the self-normalizing property of SELU activations. It applies a dropout mask during training, scaled to ensure the mean and variance of activations remain stable. The scaling factor is computed based on the dropout rate and the SELU parameters ($\lambda$ and $\alpha$). Alpha-Dropout is mathematically defined as:
[…]
@@
where, $x$ is the input activation, $\mu(x)$, $std(x)$ are mean and standard deviation of the input activation, respectively, and $mask$ is a binary mask generated with the specified dropout rate.
@@
Together, SELU activations and Alpha-Dropout ensure that SeNMo blocks maintain stable mean and variance across network layers, facilitating more reliable training and better generalization performance. Additionally, these mechanisms help mitigate training instabilities related to vanishing or exploding gradients in feedforward networks. Our network architecture consists of seven fully connected hidden layers, each followed by SELU activation and Alpha-Dropout, as illustrated in Figure 4. The number of neurons in each block is shown in the inset of Figure 4. The final fully connected layer is used to learn a latent representation of each sample, termed as the patient embedding $\textbf{x}\in\mathbb{R}^{48}$.
@@ 2.4 Training and Evaluation
2.4.1 Data Splits
For the OS task, the pan-cancer data was randomly divided into the training-validation set ($80\%$) and the hold-out test set ($20\%$) for each cancer type. The pan-cancer training was carried out by combining the training-validation cohort of all $33$ cancer types and adopting the 10-fold cross-validation with the $80-20\%$ division of samples. The training-validation cohort has $11,050$ patients, each having $\mathbb{R}^{80,697}$ features, comprising the six multi-omics modalities, gene expression, DNA methylation, miRNA expression, protein expression, DNA mutation, and the four clinical features (age, gender, race, stage). The SeNMo encoder model was trained on the training-validation cohort for the regression task of predicting the OS. C-Index was used as the evaluation metric of the hazard score predicted by the model. We used weights and biases to find the optimal set of hyperparameters for our deep learning model [107]. For the evaluation/testing of the trained model, the inference data was created by combining the held-out test set from all $33$ cancer types, resulting in $2,754$ patients, each having $\mathbb{R}^{80,697}$ features. We further tested the optimal hyperparameters of our trained model to train different combinations of the pan-cancer data modalities. We call these 1-modal, 3-modal (gene expression, DNA methylation, miRNA expression), 4-modal ($3+$protein expression), 5-modal ($4+$DNA mutation), and 6-modal (all modalities) cohorts. Although our initial model was trained on all 6 modalities, these experiments aim to see how the model performs on each of these pan-cancer cohorts where one or more of the data modalities is missing.
@@ 2.4.2 Evaluation
We evaluate SeNMo’s performance with the quantitative and statistical metrics common for survival outcome prediction and classification. For survival analysis, we evaluated the model using the C-index. For the primary cancer type classification, we generate the classification report comprising average accuracy, average precision, recall, F1-score, confusion matrix, and scatter plot. For the TLS Ratio, we employed Huber Loss. We utilized the log-rank test to determine if the survival predictions were statistically significantly different. Below, we explain the loss, evaluation metrics, and statistical tests in detail.
@@
    Loss Function: The loss being used for backpropagation in the model is a combination of three components: Cox loss, cross-entropy loss, and regularization loss. This combined loss function aims to simultaneously optimize the model’s ability to predict survival outcomes (Cox loss), encourage model-simplicity or sparsity (regularization loss), and model the likelihood of cancer types (cross-entropy loss). The overall loss is a weighted sum of these three components, where each component is multiplied by a corresponding regularization hyperparameter ($\lambda_{c}$, $\lambda_{ce}$, $\lambda_{r}$). This weighted sum allows for balancing the influence of each loss component on the optimization process. Mathematically, the overall loss can be expressed as:
@@
    • Cox proportional hazards loss ($L_{cox}$): Cox loss is a measure of dissimilarity between the predicted hazard scores and the true event times in survival analysis. It is calculated using the Cox proportional hazards model and penalizes deviations between predicted and observed survival outcomes of all individuals who are at risk at time $t_{i}$, weighted by the censoring indicator [108]. The function takes a vector of survival times for each individual in the batch, censoring status for each individual (1 if the event occurred, 0 if censored), and the predicted log hazard ratio for each individual from the neural network, and returns the Cox loss for the batch, which is used to train the neural network via backpropagation. This backpropagation encourages the model to assign higher hazards to high-risk individuals and lower predicted hazards to censored individuals or those who experience the event later. Mathematically, the Cox loss is expressed as:
@@
      where $N$ is the batch size (number of samples), $\theta_{i}$ is the predicted hazard for sample $i$, $R_{ij}$ is the indicator function that equals 1 if the survival time of sample $j$ is greater than or equal to the survival time of sample $i$, and 0 otherwise, and $\delta_{i}$ is the censoring indicator for sample $i$, which equals 1 if the event is observed for sample $i$ and 0 otherwise.
    • Cross-entropy loss ($L_{ce}$): The cross-entropy loss is a common loss function used for multi-class classification problems, particularly when each sample belongs to one of the $C$ classes. When combined with a LogSoftmax layer, the function measures how well a model’s predicted log probabilities match the true distribution across various classes. For a multi-class classification problem having $C$ classes, the model’s outputs (raw class scores or logits) are transformed into log probabilities using a LogSoftmax layer. The cross-entropy loss compares these log probabilities to the true distribution, which is usually represented in a one-hot encoded format. The loss is calculated by negating the log probability of the true class across all samples in a batch and then averaging these values. For the given output of LogSoftmax, $\log(p_{n,c})$ for each class $c$ in each sample $n$, the cross-entropy loss for a multi-class problem can be defined as:
@@
      where $N$ is the total number of samples, $C$ are the total classes, and $y_{n,c}$ is the target label for sample $n$ and class $c$, typically 1 for the true class and 0 otherwise.
@@
    Concordance Index (C-index): The C-index is a frequently used evaluation metric in survival analysis to assess the predictive accuracy of a model for the time-to-event outcomes [10]. It measures the degree to which the model’s predictions correlate with the actual survival times observed in the data. It quantifies the model’s ability to correctly rank pairs of subjects based on their predicted survival times. The C-index evaluates the probability that, in a randomly selected pair of individuals, the one who experienced the event (like death or failure) first also had a higher risk score predicted by the model. Risk score is the output of the survival model and represents the expected order of the events; the higher the score, the higher the risk of experiencing the event sooner [10]. We used the $concordance\_index$ Lifelines function to calculate the C-index [109]. This function takes the predicted hazard scores for each individual, the true event indicator (e.g., 1 if an event occurred, 0 if censored) for each individual, and the survival times (time to event or censoring) for each individual. The C-index function computes the fraction of all pairs of subjects whose predicted event times are correctly ordered among all pairs where one subject experienced an event and the other did not. C-index ranges between 0 and 1 where 0.5 is the expected result from random predictions, 1.0 is a perfect concordance, and 0.0 is perfect anti-concordance [10]. Mathematically,
@@
    where concordant pairs are pairs of individuals where the predicted survival times are correctly ordered relative to the observed survival times, tied pairs are the number of pairs where the predictions are equal or survival times are the same. Total number of evaluable pairs are the total pairs considered, excluding pairs with censoring issues or other exclusions, $\hat{S}_{i}$ and $\hat{S}_{j}$ represent the predicted risks or survival probabilities for individuals $i$ and $j$, respectively. $T_{i}<T_{j}$ implies that individual $i$ experienced the event before individual $j$, and $\delta_{i}=1$ indicates that the event for individual $i$ was observed (not censored).
@@
    Cox log-rank function: The Cox log-rank function calculates the p-value using the log-rank test based on predicted hazard scores, censor values, and the true OS times. The log-rank test is a statistical method to compare the survival distributions of two groups or more groups, where the null hypothesis is that there is no difference between the groups. It is commonly used in survival analysis to compare the observed number of events in each group to the number of events expected under the null hypothesis. For the hazard ratio $h_{i}(t)$ of group $i$ at time $t$, the hypotheses are given by,
    $\begin{split}&H_{0}:h_{1}(t)=h_{2}(t)\\ &H_{A}:h_{1}(t)=\delta h_{2}(t),\;\;\delta\neq 1\end{split}$ (11)
@@
    where $O_{i}$ is the observed number of events at time point $i$ in the sample, $E_{i}$ is the expected number of events at time point $i$ under the null hypothesis, and $N$ is the total number of observed time points.
  4. 4.
    Huber Loss: For TLS ratio prediction, we used Huber Loss, a loss function commonly used in regression tasks, known for combining the advantages of both the Mean Absolute Error (MAE) and the Mean Squared Error (MSE). It behaves differently based on the magnitude of the error; it is quadratic for small errors and linear for large errors. This characteristic makes it less sensitive to outliers than MSE and more sensitive to small errors than MAE. Huber loss function is defined as follows:
    $L_{n}=\begin{cases}0.5(y_{n}-\hat{y_{n}})^{2},&\text{if }|y_{n}-\hat{y_{n}}|\leq\delta,\\ \delta*\left(|y_{n}-\hat{y_{n}}|-0.5*\delta\right),&\text{otherwise}.\end{cases}$ (13)
    where $(y-\hat{y})$ represents the residual, which is the difference between the actual value and the predicted value, and $\delta$ is a positive threshold parameter that determines the point at which the loss function transitions from quadratic to linear behavior [110].
  5. 5.
    Wilcoxon Signed-Rank Test: To assess the agreement between the manually annotated TLS ratio and the model’s predictions, we used the Wilcoxon Signed-Rank test. This non-parametric statistical method evaluates whether there is a significant difference between the paired values, taking into account both the magnitude and direction of the differences. The null hypothesis assumes that the two distributions are statistically similar. A two-sided p-value of less than 0.05 was considered evidence of a significant difference between the two sets of ratios.
Table 3: Hyperparameters search for training.
Hyperparams Training (range)
Learning Rate [1e-6, 1e-1]
Weight Decay [1e-6, 1e-1]
Dropout [0.1, 0.65]
Batch Size [64, 128, 256, 512]
Epochs [50, 100]
Hidden Layers [1, 2, 3, 4, 5, 6, 7, 8, 9]
Hidden Neurons [2048, 1024, 512, 256, 128, 48, 32]
Optimizer [adam, sgd, rmsprop, adamw]
Learning Rate Policy [linear, exp, step, plateau, cosine]
@@
2.4.3 Hyperparameters Search
Hyperparameters are non-learnable parameters of a deep learning model and are crucial as they govern the learning process and model architecture. Hyperparameter tuning involves selecting the optimal combination of parameters that results in the best model performance. Common hyperparameters include learning rate and policy, batch size, number of epochs, weight decay, dropout type and probability, and architecture specifics such as the number of hidden layers and neurons in each layer. Methods for hyperparameter search range from grid search, where all possible combinations of parameters are evaluated; to random search, which randomly samples parameter combinations within predefined bounds. More sophisticated techniques like Bayesian optimization or using automated machine learning (AutoML) tools can dynamically adjust parameters based on previous results to find the best solutions more efficiently. We employed weights and biases [107] utility to carry out random and Bayesian methods of hyperparameters search. The list of hyperparameters we searched for training is given in Table 3. For model training, we conducted around 400 simulations to find the current hyperparameters. To further verify the performance of our model, we evaluated the model with the off-the-shelf datasets CPTAC-LSCC [57] and Moffitt’s LSCC [58]. The plot for these simulations is given in Figure 11.
2.4.4 Frameworks, Compute resources, and wall-clock times
We trained SeNMo model using the Moffitt Cancer Center’s HPC machine using one Tesla V100 32GB GPU running Ubuntu 22.04.4 and CUDA 12.2. The entire code was developed in Python and PyTorch frameworks. The software frameworks and corresponding packages used in our codebase are given in Table 4. Training time for our current 83.33 Million parameter SeNMo encoder is approximately 11 hours. We conducted the hyperparameters search of the pan-cancer model for approximately 20 days using multiple GPUs in parallel. Finetuning the trained model on a given data having around 150 patients approximately takes 15 minutes.
Figure 5: Study design and simulations structure for SeNMo across different learning regimes, datasets, and tasks. The baseline model (row 1) is first trained on TCGA data with 33 cancer types for overall survival (OS) prediction. Rows 2–5 represent variations: red-bordered boxes indicate a change from the baseline (e.g., out-of-distribution task and/or data), while green-bordered boxes align with the baseline. Simulations include OS prediction on both seen and unseen data (rows 2 and 3) and new tasks such as primary cancer type classification and TLS ratio prediction on seen and unseen datasets (rows 4 and 5).
2.5 Study Design
An overview of the various simulations conducted to evaluate the capabilities of the SeNMo model across different learning regimes, tasks, and datasets is shown in Figure 5. The study design included multiple learning regimes, each designed to assess the model’s adaptability, generalizability, and robustness. The baseline model was initially trained on TCGA dataset comprising 33 different cancer types for OS prediction. The subsequent learning regimes explored different data variations and tasks, which we call out-of-distribution simulations because the model had not encountered such data/task in baseline learning. These scenarios included OS prediction on both seen and unseen datasets, as well as tasks such as primary cancer type classification on seen data and TLS ratio prediction on unseen data.
Figure 6: Pan-cancer C-Index results for OS prediction. The SeNMo model was trained and evaluated using different combinations of data modalities. Training and validation were carried out on the $80\%$ of the total data, whereas inference was done on the $20\%$ held-out test set. As the number of modalities increased in the pan-cancer data, the model’s performance improved, as depicted by the upward trend of C-Index. All the results shown here are statistically significant, i.e., $p<0.05$.